Simplify:
11(1/n - 1) - 9/n if n=1/101.
11(1/1/101 - 1) - 1/1/101.
We are stuck with what to do with the fractions .... I recall that 1/1/101 = 1(101) but can not remember what property to use.
Help!

Answer by stanbon(75887)

Simplify:
11(1/n - 1) - 9/n if n=1/101.

--------------
If I assume 1/n-1 means (1/n)-1 I would get the following results:
11((1/n)-1) - (9/n)
= (11/n) - 11 - (9/n)
= (2/n) - 11
Now let n = 1/101
= 2/(1/101) - 11
= 202 - 11
= 191
=============
Cheers,
Stan H.